OpenStudy (anonymous):

If we toss one die 1000 times, how many times would we expect to get a 5? Answer about 56 about 100 about 168 about 200 1 points Question 8 Three coins are tossed during an experiment. It is equally as likely that no heads will appear as it is that exactly one head will appear. Answer True False

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

u dont have to answer #2 i got that one its false

OpenStudy (anonymous):

about 200

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

yep 1000/5=200

OpenStudy (anonymous):

ty! can u help me with a few more?

OpenStudy (anonymous):

yep

OpenStudy (mayankdevnani):

lolz...i don't think so

OpenStudy (mayankdevnani):

@ganeshie8 help!

OpenStudy (mayankdevnani):

@sidsiddhartha

OpenStudy (mayankdevnani):

@wio @Kainui

OpenStudy (anonymous):

If we draw a card from a deck of 52 playing cards, what is the probability that it is a queen? Answer 1/52 1/26 1/13 1/4

OpenStudy (mayankdevnani):

there are 4 queens.= `favourable outcomes` total outcomes =52

OpenStudy (mayankdevnani):

\[\large \bf probability=\frac{4}{52}=\frac{1}{13}\]

OpenStudy (anonymous):

so the answer is (A) 1/52?

OpenStudy (mayankdevnani):

no !! there are 4 queens

OpenStudy (anonymous):

oh ok opps....

OpenStudy (anonymous):

In a deck of 52 playing cards, there are 4 aces. What is the probability that if you cut the deck, you’ll get an ace? Answer 1/18 1/12 1/14 1/13

OpenStudy (anonymous):

wait thats the wrong answer key the right one is a. 1/8 b. 1/12 c. 1/4 d. 1/13

OpenStudy (anonymous):

@ganeshie8

OpenStudy (only4u):

i think the answer of 1st qn is 168